1. The control panel on a spaceship contains a light that blinks every 3.01 s as observed by an astronaut in the ship. If the spaceship is moving past Earth with a speed of 0.743c, determine the proper time interval between blinks and the time interval between blinks as observed by a person on Earth.

A. the proper time interval (in s) between blinks

___________ s

B. the time interval (in s) between blinks as observed by a person on Earth

____________ s

2. An astronaut at rest on Earth has a heartbeat rate of 88 beats/min. When the astronaut is traveling in a spaceship at 0.83c, what will this rate be as measured by an observer also in the ship and an observer at rest on Earth?

A. an observer also in the ship
____________beats/min

B. an observer at rest on Earth
____________beats/min

Can you please help me with both questions? I'll be sure to rate your response. Thank you!!
Answer #1

1. Given,

The time interval is taken by the light to blink, t = 3.01 sec

The speed, v = 0.743 c

A.

The proper time interval between the blinks, tp = 3.01 sec

B.

The time interval between blinks as observed by a person on Earth is, t = tp / √[1-(v/c)2]

t = 3.01 / √[1 - (0.743c/c)2]

t = 3.01 / √(1 - 0.552)

= 4.49 sec